OpenStudy (anonymous):

Find exact solutions 6x^2-36=0.

OpenStudy (amoodarya):

6x^2-36=0. /6 x^2-6=0 x^2=6

OpenStudy (agent0smith):

You can also factor it, after dividing both sides by 6. It's a difference of two squares, x^2-6=0